OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (1, −3) and has a slope of 2? 2x − y = 5 2x − y = −5 2x + y = 5 2x + y = −5

OpenStudy (texaschic101):

y = mx + b slope(m) = 2 (1,-3)...x = 1 and y = -3 now we sub -3 = 2(1) + b -3 = 2 + b -3 - 2 = b -5 = b so our equation is : y = 2x - 5 But we need it in standard form : Ax + By = C y = 2x - 5 -- subtract 2x from both sides -2x + y = 5 -- multiply by -1 to make x positive 2x - y = -5 (standard form)

OpenStudy (anonymous):

so the answer is 2x-y=-5?

OpenStudy (texaschic101):

yep...thats the answer

OpenStudy (anonymous):

thank you soo much

OpenStudy (texaschic101):

not a problem...happy to help :)

OpenStudy (anonymous):

this is my last question please help

OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (4, 0) and (2, 6)? y = 4x − 2 y = 2x − 4 y = −3(x − 4) y = 3(x + 4)